# Solvable Tuple Patterns and Their Applications to Program Verification

### v1

@@ Appendix A Proofs
A.1 Proofs of Soundness, Completeness, and Minimality
This section proves the main properties of our algorithm: soundness, completeness, minimality, the sufficiency of polynomial size data, and complexity results. The proofs in this section are for the STPs and inference algorithm as defined in Sections 2.1 and 2.2; the extensions in Section 2.3 are treated in Section A.2.
We first show basic properties in Sections A.1.1A.1.3. The proofs of the main properties use only those properties; therefore, if we extend the rewriting relation and the pattern language (as we did so in Section 2.3), it is sufficient to reprove those properties. The basic properties required for each main property are summarized in Table 3.
@@ A.1 Proofs of Soundness, Completeness, and Minimality
Soundness (Theorem 2.1) Lemmas A.1 and A.2
Soundness 2 (Theorem 2.3) Lemmas A.1, A.3, and A.9
Completeness (Theorem 2.4) Lemmas A.1 and A.10
Minimality (Theorem 2.5) Lemmas A.1, A.3, A.5, A.6,A.9, A.10, and A.11
Required Data Size (Theorem 2.6) Lemma A.18, in addition to all the lemmas above
Complexity of Decision Problems (Theorem 2.8) Lemmas A.4, A.5, and A.6
A.1.1 Properties of $\longrightarrow$.
@@
Let $\widetilde{x}$ be mutually distinct variables. If $(t,[M/\widetilde{x}])\longrightarrow^{n}(t^{\prime},\Theta^{\prime})$, then there exists $t^{\prime\prime}$ such that $(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{n}(t^{\prime\prime},\Theta^{\prime})$ and $t^{\prime}=[t^{\prime\prime}/\widetilde{x}]t$.
@@
    If $t_{1}\leadsto t_{1}^{\prime}$ and $\mathcal{L}(t_{0})\subseteq\mathcal{L}(t_{1})$, then $\langle t_{1}\leadsto t_{1}^{\prime}\rangle(t_{0})$ is well defined.
@@
    If $\langle t_{1}\leadsto t_{1}^{\prime}\rangle(t_{0})$ is well defined, then $\mathcal{L}(t_{0})\subseteq\mathcal{L}(t_{1})$ if and only if $\mathcal{L}(\langle t_{1}\leadsto t_{1}^{\prime}\rangle(t_{0}))\subseteq\mathcal{L}(t_{1}^{\prime})$; and $\mathcal{L}(t_{0})\supseteq\mathcal{L}(t_{1})$ if and only if $\mathcal{L}(\langle t_{1}\leadsto t_{1}^{\prime}\rangle(t_{0}))\supseteq\mathcal{L}(t_{1}^{\prime})$.
@@
Case analysis on the rule used for deriving $t_{1}\leadsto t_{1}^{\prime}$. Since the other cases are similar or simpler, we discuss only the case for PR-Prefix. We may assume, without loss of generality, $t_{1}=(p_{2}p_{1},p_{2},\ldots,p_{n})$ and $t_{1}^{\prime}=(p_{1},p_{2},\ldots,p_{n})$ with $p_{2}\neq\epsilon$. To show (1), assume $\mathcal{L}(t_{0})\subseteq\mathcal{L}(t_{1})$. Then for any $(s_{1},s_{2},\ldots,s_{n})\in\mathcal{L}(t_{0})$, $s_{2}$ must be a prefix of $s_{1}$. Thus $t_{0}$ must also be of the form $(q_{2}q_{1},q_{2},\ldots,q_{n})$, and hence $\langle t_{1}\leadsto t_{1}^{\prime}\rangle(t_{0})$ is well defined.
@@
$\mathcal{L}(t_{1}^{\prime})$ $=\{(w_{2}\backslash w_{1},w_{2},\ldots,w_{n})\mid(w_{1},\ldots,w_{n})\in\mathcal{L}(t_{1})\}$
$\mathcal{L}(t_{1})$ $=\{(w_{2}w_{1},w_{2},\ldots,w_{n})\mid(w_{1},\ldots,w_{n})\in\mathcal{L}(t^{\prime}_{1})\},$
@@
It is sufficient to consider the cases where $t_{1}\neq t_{2}$ and the parts involved in the two reductions overlap. Without loss of generality, we may assume that the first element $p_{1}$ is principal in $t\leadsto t_{1}$. We thus need to consider only the following cases.
@@ A.1.3 Correspondence between $\longrightarrow$ and $\leadsto$
  • Case PR-Epsilon: We may assume, without loss of generality, $t=(\epsilon,p_{2},\ldots,p_{n})$ and $t^{\prime}=(p_{2},\ldots,p_{n})$. The required result holds for $t^{\prime\prime}=(\epsilon,x_{2},\ldots,x_{n})$, $M^{\prime}=M\mathbin{\uparrow_{j}}$ and $\widetilde{y}=(x_{2},\ldots,x_{n})$.
Lemma A.11.
[…]
Proof.
[…]
  • Case R-Prefix: Without loss of generality, we can assume:
    $\widetilde{x}=(x_{1},\ldots,x_{n})\qquad\widetilde{y}=(x_{1}^{\prime},x_{2},\ldots,x_{n})\qquad t=(x_{2}x_{1}^{\prime},x_{2},\ldots,x_{n})$
    $t^{\prime}=(p_{1},\ldots,p_{n})\qquad[t^{\prime}/\widetilde{y}]t=(p_{2}p_{1},p_{2},\ldots,p_{n})$
    $M[*][1]=M[*][2]\cdot M^{\prime}[*][1]\qquad M[*][j]=M^{\prime}[*][j]\mbox{ for $j=2,\ldots,n$}.$
[…]
  • Case R-CPrefix: Similar to the case above.
  • Case R-Epsilon: Without loss of generality, we can assume:
    $\widetilde{x}=(x_{1},\ldots,x_{n})\qquad\widetilde{y}=(x_{2},\ldots,x_{n})\qquad t=(\epsilon,x_{2},\ldots,x_{n})$
    $t^{\prime}=(p_{2},\ldots,p_{n})\qquad[t^{\prime}/\widetilde{y}]t=(\epsilon,p_{2},\ldots,p_{n})\qquad M^{\prime}=M\mathbin{\uparrow_{j}}.$
    By the assumption $\mathcal{L}(t_{0})\subseteq\mathcal{L}([t^{\prime}/\widetilde{y}]t)$, $t_{0}$ must be of the form $(\epsilon,q_{2},\ldots,q_{n})$. Let $t_{0}^{\prime}=(q_{2},\ldots,q_{n})$. Then, we have $t_{0}\leadsto t_{0}^{\prime}$, $\mathcal{L}(t_{0}^{\prime})\subseteq\mathcal{L}(t^{\prime})$, and $M^{\prime}\models t_{0}^{\prime}$. Furthermore, $\mathcal{L}(t_{0}^{\prime})=\mathcal{L}(t^{\prime})$ implies $\mathcal{L}(t_{0})=\mathcal{L}([t^{\prime}/\widetilde{y}]t)$ as required.
A.1.4 Proof of Soundness (Theorems 2.1 and 2.3)
We are now ready to prove the soundness theorem.
@@
A.1.5 Proof of Completeness (Theorem 2.4).
@@
for some $\Theta$. By Lemma A.1, we have $(t_{1},[M^{\prime}/\widetilde{y}])\longrightarrow^{*}([t^{\prime}/\widetilde{y}]t_{1},\Theta)=(t,\Theta)$. Thus, we have $(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{*}(t,\Theta)$ as required. ∎
A.1.6 Proof of Minimality (Theorem 2.5).
@@
In the induction step, we have $(\widetilde{x},[M/\widetilde{x}])\longrightarrow(t^{\prime},[M^{\prime}/\widetilde{y}])\longrightarrow^{k}(t_{1},\Theta_{1})$. By Lemma A.1, there exists $t_{2}$ such that $(\widetilde{y},[M^{\prime}/\widetilde{y}])\longrightarrow^{k}(t_{2},\Theta_{1})$ and $t_{1}=[t_{2}/\widetilde{y}]t^{\prime}$. By Lemma A.11, there exists $t_{0}^{\prime}$ such that $t_{0}\leadsto^{*}t_{0}^{\prime}$ and $\mathcal{L}(t_{0}^{\prime})\subseteq\mathcal{L}(t_{2})$ with $M^{\prime}\models t_{0}^{\prime}$. Furthermore, $\mathcal{L}(t_{0}^{\prime})=\mathcal{L}(t_{2})$ implies $\mathcal{L}(t_{0})=\mathcal{L}([t_{2}/\widetilde{y}]t^{\prime})=\mathcal{L}(t_{1})$. Thus, we can apply the induction hypothesis to obtain $\mathcal{L}(t_{0}^{\prime})=\mathcal{L}(t_{2})$, which implies $\mathcal{L}(t_{0})=\mathcal{L}(t_{1})$ as required. ∎
A.1.7 Alternative Proof of Minimality
We first extend $\langle t\leadsto t^{\prime}\rangle$ to a partial operation on data.
Definition A.2.
Suppose $t\leadsto t^{\prime}$. We define $\langle t\leadsto t^{\prime}\rangle(M)$ by case analysis on the rule used for deriving $t\leadsto t^{\prime}$, as follows.
  • Case PR-Prefix: In this case, $t=(p_{1},\ldots,p_{n})$ and $t^{\prime}=(p^{\prime}_{1},\ldots,p^{\prime}_{n})$ with $p_{j}=p_{i}\cdot p^{\prime}_{j}$ and $p_{i}\neq\epsilon$ for some $i,j$ and $p^{\prime}_{k}=p_{k}$ for $k\neq j$. If $\mathit{shape}(M)=(m,n)$ and $M[*][j]=M[*][i]\cdot\widetilde{s}$, then $\langle t\leadsto t^{\prime}\rangle(M):=M\{j\mapsto\widetilde{s}\}$. Otherwise, $\langle t\leadsto t^{\prime}\rangle(M)$ is undefined.
  • Case PR-CPrefix: In this case, $t=(p_{1},\ldots,p_{n})$ and $t^{\prime}=(p^{\prime}_{1},\ldots,p^{\prime}_{n})$ with $p_{j}=ap^{\prime}_{j}$ for some $j$ and $p^{\prime}_{k}=p_{k}$ for $k\neq j$. If $\mathit{shape}(M)=(m,n)$ and $M[*][j]=a\cdot\widetilde{s}$, then $\langle t\leadsto t^{\prime}\rangle(M):=M\{j\mapsto\widetilde{s}\}$. Otherwise, $\langle t\leadsto t^{\prime}\rangle(M)$ is undefined.
  • Case PR-Epsilon: In this case, $t=(p_{1},\ldots,p_{j-1},\epsilon,p_{j+1},\ldots,p_{n})$ and $t^{\prime}=(p_{1},\ldots,p_{j-1},p_{j+1},\ldots,p_{n})$. If $\mathit{shape}(M)=(m,n)$ and $M[*][j]=\widetilde{\epsilon}$, then $\langle t\leadsto t^{\prime}\rangle(M):=M\mathbin{\uparrow_{j}}$. Otherwise, $\langle t\leadsto t^{\prime}\rangle(M)$ is undefined.
Lemma A.14.
If $t\leadsto t^{\prime}$ and $M,\Theta\models t$, then $\langle t\leadsto t^{\prime}\rangle(M)$ is well defined and $\langle t\leadsto t^{\prime}\rangle(M),\Theta\models t^{\prime}$.
Proof.
This follows immediately from the definition. ∎
Lemma A.15 (stronger version of Lemma A.15).
[…]
Proof.
[…]
Lemma A.16.
If $(\widetilde{x},[M/\widetilde{x}])\longrightarrow(t_{1},[M_{1}/\widetilde{y}])\longrightarrow^{k}(t,\Theta)$, then there exists $t^{\prime}$ such that $(\widetilde{y},[M_{1}/\widetilde{y}])\longrightarrow^{k}(t^{\prime},\Theta)$ with $t=[t^{\prime}/\widetilde{y}]t_{1}$. Furthermore, $\langle t_{1}\leadsto\widetilde{y}\rangle=\langle[t^{\prime}/\widetilde{y}]t_{1}\leadsto t^{\prime}\rangle$ holds.
Proof.
[…]
Lemma A.17.
If $M\models t$ and $\langle t_{1}\leadsto t_{2}\rangle(t)$ is well defined, then $\langle t_{1}\leadsto t_{2}\rangle(M)\models\langle t_{1}\leadsto t_{2}\rangle(t)$.
Proof.
This follows by a straightforward case analysis of $t_{1}\leadsto t_{2}$. ∎
Proof of Theorem 2.5.
In the induction step, we have $(\widetilde{x},[M/\widetilde{x}])\longrightarrow(t^{\prime},[M^{\prime}/\widetilde{y}])\longrightarrow^{k}(t_{1},\Theta_{1})$. By Lemma A.16, there exists $t_{2}$ such that $(\widetilde{y},[M^{\prime}/\widetilde{y}])\longrightarrow^{k}(t_{2},\Theta_{1})$ and $t_{1}=[t_{2}/\widetilde{y}]t^{\prime}$, with $\langle t^{\prime}\leadsto\widetilde{y}\rangle=\langle t_{1}\leadsto t_{2}\rangle$. By the assumption $\mathcal{L}(t_{0})\subseteq\mathcal{L}(t_{1})$ and Lemma A.4, we have:
$\langle t_{1}\leadsto t_{2}\rangle(t_{0})\mbox{ is well-defined.}$
$\mathcal{L}(\langle t_{1}\leadsto t_{2}\rangle(t_{0}))\subseteq\mathcal{L}(t_{2})$
$\mathcal{L}(\langle t_{1}\leadsto t_{2}\rangle(t_{0}))\supseteq\mathcal{L}(t_{2})\mbox{ implies }\mathcal{L}(t_{0})\supseteq\mathcal{L}(t_{1}).$
By Lemmas A.15 and A.17 with the assumption $M\models t_{0}$, we have
$M^{\prime}=\langle t^{\prime}\leadsto\widetilde{y}\rangle(M)\models\langle t^{\prime}\leadsto\widetilde{y}\rangle(t_{0})=\langle t_{1}\leadsto t_{2}\rangle(t_{0}).$
Thus, we can apply the induction hypothesis to $(\widetilde{y},[M^{\prime}/\widetilde{y}])\longrightarrow^{k}(t_{2},\Theta_{1})$ and obtain $\mathcal{L}(\langle t_{1}\leadsto t_{2}\rangle(t_{0}))=\mathcal{L}(t_{2})$. Thus, we have $\mathcal{L}(t_{0})=\mathcal{L}(t_{1})$ as required. ∎
A.1.8 Proof of Other Properties.
@@ A.1.7 Alternative Proof of Minimality
To prove Theorem 2.6, we first define canonical learning data $M^{t}$ for each solvable tuple pattern $t$. Let $t=(p_{1},\ldots,p_{n})$ be a solvable tuple pattern. Since $t$ is solvable, $t$ contains at most $n$ distinct variables $x_{1},\ldots,x_{k}$ with $k\leq n$. (Note that, at each reduction step by $\leadsto$, the set of variables occurring in the tuple pattern does not change, and the number of elements may only monotonically decrease.) Let $a,b$ be two distinct variables of $\Sigma$, and $\alpha_{1},\ldots,\alpha_{n}$ be binary prefix codewords for $1,\ldots,n$ using $a,b$ such that $|\alpha_{i}|=O(\log n)$. Let $\beta_{i}\,(i=1,\ldots,n)$ be $a/b,b/a]\alpha_{i}$. We define $M^{t}$ as: [^14] The canonical data $M^{t}$ depends on the choice of $a,b\in\Sigma$ and binary prefix coding, but the choice does not matter for the discussion below.
@@
We use the following properties of canonical data.
@@
  • Case PR-Const: Similar to the above case.
  • Case PR-Epsilon: We may assume without loss of generality that $t_{1}=(\epsilon,q_{2},\ldots,q_{n})$ and $t_{1}^{\prime}=(q_{2},\ldots,q_{n})$. By the assumption $M^{t},\Theta\models t_{1}$, $M^{t}[*][1]=\widetilde{\epsilon}$. Therefore, $t$ must be of the form $(\epsilon,p_{2},\ldots,p_{n})$, with $M^{t^{\prime}}\models t_{1}^{\prime}$ for $t^{\prime}=(p_{2},\ldots,p_{n})$. By the induction hypothesis, we have $\mathcal{L}(t^{\prime})\subseteq\mathcal{L}(t_{1}^{\prime})$. Thus, we have $\mathcal{L}(t)\subseteq\mathcal{L}(t_{1})$ as required.

### v2

@@ Appendix A Proofs
A.1 Proofs and Additional Definitions for Sections 2 and 3
This section provides proofs omitted in Sections 2 and 3.
We first show basic properties in Sections A.1.1A.1.3. The dependency of each main property on basic properties is summarized in Table 3, so that it is easy to discuss properties of the extensions and variations of STPs considered in Section 4.
@@ A.1 Proofs and Additional Definitions for Sections 2 and 3
Soundness (Theorem 2.1) Lemma A.2
Soundness 2 (Theorem 3.1) Lemmas A.1, A.3, and A.10
Completeness (Theorems 3.2 an 3.3) Lemmas A.1 and A.11 and Theorem 2.1
Minimality (Theorem 3.4) Lemmas A.1, A.3, A.4, A.6, A.7, Lemma A.8, and A.10
Characteristic Data (Theorem 3.5) Lemmas A.14 and A.15, in addition to all the lemmas above
Learnability of STPs (Theorems 3.6 and 3.8) Soundness and the existence of characteristic data (Theorems 2.1 and  3.5)
Learnability of CSTPs (Theorems 3.6 and 3.7) Soundness, completeness and Lemma A.16
Complexity of Decision Problems (Theorem 3.9) Lemmas A.4, A.6, and A.7
Decidability of Satisfiability (Theorem 3.10) Lemma A.4 and decidability of word equations
A.1.1 Properties of $\longrightarrow$
@@
Let $\widetilde{x}$ be mutually distinct variables. If $(t,[M/\widetilde{x}])\longrightarrow^{n}(t^{\prime},\Theta^{\prime})\;\mathbin{\ \not\!\!\longrightarrow}$, then there exists $t^{\prime\prime}$ such that $(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{n}(t^{\prime\prime},\Theta^{\prime})\;\mathbin{\ \not\!\!\longrightarrow}$ and $t^{\prime}=[t^{\prime\prime}/\widetilde{x}]t$.
@@ A.1.2 Properties of $\leadsto$
  • Case PR-Suffix: In this case, $t_{1}=(p_{1},\ldots,p_{n})$ and $t_{2}=(p^{\prime}_{1},\ldots,p^{\prime}_{n})$ with $p_{j}=p^{\prime}_{j}\cdot p_{i}$ and $p_{i}\neq\epsilon$ for some $i,j$ and $p^{\prime}_{k}=p_{k}$ for $k\neq j$. If $t_{0}$ is of the form $(q_{1},\ldots,q_{n})$ and $q_{i}$ is a suffix of $q_{j}$, then $\langle t_{1}\leadsto t_{2}\rangle(t_{0}):=(q^{\prime}_{1},\ldots,q^{\prime}_{n})$ where $q_{j}^{\prime}=q_{j}/q_{i}$ and $q_{k}^{\prime}=q_{k}$ for $k\neq j$. Otherwise, $\langle t_{1}\leadsto t_{2}\rangle(t_{0})$ is undefined.
@@
  • Case PR-CSuffix: In this case, $t_{1}=(p_{1},\ldots,p_{n})$ and $t_{2}=(p^{\prime}_{1},\ldots,p^{\prime}_{n})$ with $p_{j}=p^{\prime}_{j}a$ for some $j$ and $p^{\prime}_{k}=p_{k}$ for $k\neq j$. If $t_{0}$ is of the form $(q_{1},\ldots,q_{n})$ and $a$ is a suffix of $q_{j}$, then $\langle t_{1}\leadsto t_{2}\rangle(t_{0}):=(q^{\prime}_{1},\ldots,q^{\prime}_{n})$ where $q_{j}^{\prime}=q_{j}/a$ and $q_{k}^{\prime}=q_{k}$ for $k\neq j$. Otherwise, $\langle t_{1}\leadsto t_{2}\rangle(t_{0})$ is undefined.
@@
    If $t_{1}\leadsto t_{1}^{\prime}$ and $\mathcal{L}(t_{0})\subseteq\mathcal{L}([t/\widetilde{x}]t_{1})$, then $\langle t_{1}\leadsto t_{1}^{\prime}\rangle(t_{0})$ is well defined and $t_{0}\leadsto^{*}\langle t_{1}\leadsto t_{1}^{\prime}\rangle(t_{0})$.
@@
    If $\langle t_{1}\leadsto t_{1}^{\prime}\rangle(t_{0})$ is well defined, then $\mathcal{L}(t_{0})\subseteq\mathcal{L}([t/\widetilde{x}]t_{1})$ if and only if $\mathcal{L}(\langle t_{1}\leadsto t_{1}^{\prime}\rangle(t_{0}))\subseteq\mathcal{L}([t/\widetilde{x}]t_{1}^{\prime})$; and $\mathcal{L}(t_{0})\supseteq\mathcal{L}([t/\widetilde{x}]t_{1})$ if and only if $\mathcal{L}(\langle t_{1}\leadsto t_{1}^{\prime}\rangle(t_{0}))\supseteq\mathcal{L}([t/\widetilde{x}]t_{1}^{\prime})$.
@@
Case analysis on the rule used for deriving $t_{1}\leadsto t_{1}^{\prime}$. Since the other cases are similar or simpler, we discuss only the case for PR-Prefix. We may assume, without loss of generality, $t_{1}=(p_{2}p_{1},p_{2},\ldots,p_{n})$ and $t_{1}^{\prime}=(p_{1},p_{2},\ldots,p_{n})$ with $p_{2}\neq\epsilon$. To show (1), assume $\mathcal{L}(t_{0})\subseteq\mathcal{L}([t/\widetilde{x}]t_{1})$. Then for any $(s_{1},s_{2},\ldots,s_{n})\in\mathcal{L}(t_{0})$, $s_{2}$ must be a prefix of $s_{1}$. Thus $t_{0}$ must also be of the form $(q_{2}q_{1},q_{2},\ldots,q_{n})$. (To see this, suppose $t_{0}=(q_{1}^{\prime},q_{2},\ldots)$ and $q_{2}$ is not a prefix of $q_{1}^{\prime}$. Then either (i) $q_{1}^{\prime}q_{1}^{\prime\prime}=q_{2}$ with $q_{1}^{\prime\prime}\neq\epsilon$ (i.e., $q_{1}^{\prime}$ is a strict prefix of $q_{2}$) or (ii) $q_{1}^{\prime}=q\alpha_{1}q_{1}^{\prime\prime}$ and $q_{2}=q\alpha_{2}q_{2}^{\prime\prime}$ with $\alpha_{1},\alpha_{2}\in\Sigma\cup\mathcal{V}$ and $\alpha_{1}\neq\alpha_{2}$ (i.e., $\alpha_{1},\alpha_{2}$ are the first letters or variables that differ between $q_{1}^{\prime}$ and $q_{2}$). In case (i), let $\theta$ be a substitution such that $\theta q_{1}^{\prime\prime}\neq\epsilon$, and in case (ii), let $\theta$ be a substitution such that $\theta\alpha_{1}\neq\theta\alpha_{2}$. Then $s_{2}=\theta q_{2}$ is not a prefix of $s_{1}=\theta q_{1}^{\prime}$. Thus, $q_{2}$ must be a prefix of $q_{1}^{\prime}$. ) Therefore, $\langle t_{1}\leadsto t_{1}^{\prime}\rangle(t_{0})$ is well defined, and $t_{0}\leadsto^{*}\langle t_{1}\leadsto t_{1}^{\prime}\rangle(t_{0})$.
@@
$\mathcal{L}([t/\widetilde{x}]t_{1}^{\prime})$ $=\{(w_{2}\backslash w_{1},w_{2},\ldots,w_{n})\mid(w_{1},\ldots,w_{n})\in\mathcal{L}([t/\widetilde{x}]t_{1})\}$
$\mathcal{L}([t/\widetilde{x}]t_{1})$ $=\{(w_{2}w_{1},w_{2},\ldots,w_{n})\mid(w_{1},\ldots,w_{n})\in\mathcal{L}([t/\widetilde{x}]t^{\prime}_{1})\},$
@@
Next, we will prove the weak confluence of $\leadsto$ (up to permutations). To that end, we prepare the following lemma.
Lemma A.5.
[…]
Proof.
[…]
Remark A.1.
[…]
@@
It is sufficient to consider the cases where $t_{1}\neq t_{2}$ and the parts involved in the two reductions overlap. Without loss of generality, we may assume that the first element $p_{1}$ is principal in $t\leadsto t_{1}$. We thus need to consider only the following cases; due to the symmetry between prefix and suffix, the cases involving only suffixes are omitted.
@@
[…]
  • Case where $t=(p_{2}p_{1}^{\prime},p_{2},\ldots,p_{n})\leadsto(p_{1}^{\prime},p_{2},\ldots,p_{n})=t_{1}$ and $t=(p_{1}^{\prime\prime}p_{3},p_{2},p_{3},\ldots,p_{n})\leadsto(p_{1}^{\prime\prime},p_{2},p_{3},\ldots,p_{n})=t_{2}$ (i.e., $p_{1}$ is principal in both reductions, and $p_{2}$ is auxiliary in the former reduction while $p_{3}$ is auxiliary in the latter; $t\leadsto t_{1}$ uses PR-Prefix, $t\leadsto t_{2}$ uses PR-Suffix). In this case, $p_{2}p_{1}^{\prime}=p_{1}^{\prime\prime}p_{3}$; so, either (i) $p_{1}^{\prime\prime}=p_{2}p_{1}^{\prime\prime\prime}$ and $p_{1}^{\prime}=p_{1}^{\prime\prime\prime}p_{3}$, or (ii) $p_{2}=p_{1}^{\prime\prime}p_{2}^{\prime}$ and $p_{3}=p_{2}^{\prime}p_{1}^{\prime}$. In case (i), we have:
    In case (ii), we have:
  • Case where $t=(p_{2}p_{1}^{\prime},p_{2},\ldots,p_{n})\leadsto(p_{1}^{\prime},p_{2},\ldots,p_{n})=t_{1}$ and $t=(p_{1},p_{2}^{\prime}p_{3},p_{3},\ldots,p_{n})\leadsto(p_{1},p^{\prime}_{2},p_{3},\ldots,p_{n})=t_{2}$ (i.e. the auxiliary element $p_{2}$ in the former reduction is principal in the latter, and $t\leadsto t_{1}$ uses PR-Prefix while $t\leadsto t_{2}$ uses PR-Suffix). In this case, we have $p_{1}=p_{2}p_{1}^{\prime}$ and $p_{2}=p_{2}^{\prime}p_{3}$. Then, we have:
    as required.
@@ A.1.3 Correspondence between $\longrightarrow$ and $\leadsto$
  • Cases for PR-Suffix and PR-CSuffix: analogous to the cases for PR-Prefix and PR-CPrefix respectively.
  • Case PR-Epsilon: We may assume, without loss of generality, $t=(\epsilon,p_{2},\ldots,p_{n})$ and $t^{\prime}=(p_{2},\ldots,p_{n})$. The required result holds for $t^{\prime\prime}=(\epsilon,x_{2},\ldots,x_{n})$, $M^{\prime}=M\mathbin{\uparrow_{1}}$ and $\widetilde{y}=(x_{2},\ldots,x_{n})$.
A.1.4 Proof of Soundness (Theorems 2.1 and 3.1)
Theorem 2.1 is a trivial consequence of Lemma A.2.
@@
A.1.5 Proofs of Completeness (Theorems 3.2 and 3.3)
@@
for some $\Theta$. By Lemma A.1, we have $(t_{1},[M^{\prime}/\widetilde{z}])\longrightarrow^{*}([t^{\prime}/\widetilde{z}]t_{1},\Theta)=(t,\Theta)$. Thus, we have $(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{*}(t,\Theta)$ as required. ∎
Proof of Theorem 3.3.
The condition $M\models\mathtt{CTPinf}(M)$ follows immediately from Theorem 2.1.
Suppose $\gamma=t_{1}\land\cdots\land t_{m}$ is a CSTP and $M\models\gamma$. Then there exists $\Theta_{k}$ such that $\Theta_{k}t_{k}=M$ for each $k\in\{1,\ldots,m\}$. Let $\widetilde{x}=x_{1},\ldots,x_{\ell}$ be the variables $x$ such that $\Theta_{k}(x)=\widetilde{\epsilon}$. Then we have $\Theta_{k}M=[\widetilde{\epsilon}/\widetilde{x}]t_{k}$, i.e., $M\models_{s}[\widetilde{\epsilon}/\widetilde{x}]t_{k}$. By Theorem 3.2, we have $(\widetilde{y},[M/\widetilde{y}])\longrightarrow^{*}([\widetilde{\epsilon}/\widetilde{x}]t_{k},\Theta)$. Let $t^{\prime}_{k}$ be an STP such that $([\widetilde{\epsilon}/\widetilde{x}]t_{k},\Theta)\longrightarrow^{*}(t^{\prime}_{k},\Theta^{\prime})\;\mathbin{\ \not\!\!\longrightarrow}$. Then $t^{\prime}_{k}$ is a conjunct of $\mathtt{CTPinf}(M)$ and $\mathcal{L}(t^{\prime}_{k})\subseteq\mathcal{L}([\widetilde{\epsilon}/\widetilde{x}]t_{k})$. Thus, we have $\mathcal{L}(\mathtt{CTPinf}(M))\subseteq\mathcal{L}(t^{\prime}_{k})\subseteq\mathcal{L}([\widetilde{\epsilon}/\widetilde{x}]t_{k})\subseteq\mathcal{L}(t_{k})$ for every $k\in\{1,\ldots,m\}$, which implies $\mathcal{L}(\mathtt{CTPinf}(M))\subseteq\mathcal{L}(\gamma)$. ∎
A.1.6 Proof of Minimality (Theorem 3.4)
@@
In the induction step, we have $(\widetilde{x},[M/\widetilde{x}])\longrightarrow(t^{\prime},[M^{\prime}/\widetilde{y}])\longrightarrow^{k}(t_{1},\Theta_{1})$. By Lemma A.10, we have $t^{\prime}\leadsto\widetilde{y}$. By Lemma A.1, there exists $t_{2}$ such that $(\widetilde{y},[M^{\prime}/\widetilde{y}])\longrightarrow^{k}(t_{2},\Theta_{1})$ and $t_{1}=[t_{2}/\widetilde{y}]t^{\prime}$. By the assumption $\mathcal{L}(t_{0})\subseteq\mathcal{L}(t_{1})=[t_{2}/\widetilde{y}]t^{\prime}$ and Lemma A.4 with $t_{1}\leadsto\widetilde{y}$, we have (i) $t^{\prime}_{0}:=\langle t^{\prime}\leadsto\widetilde{y}\rangle(t_{0})$ is well-defined; (ii) $\mathcal{L}(t^{\prime}_{0})\subseteq\mathcal{L}([t_{2}/\widetilde{y}]\widetilde{y})=\mathcal{L}(t_{2})$; and (iii) $\mathcal{L}(t^{\prime}_{0})\subseteq\mathcal{L}(t_{2})$ implies $\mathcal{L}(t_{0})=\mathcal{L}(t_{1})$. Furthermore, by $\mathcal{T}\ni t_{0}\leadsto^{*}t_{0}^{\prime}$ and Lemma A.13, $t_{0}^{\prime}$ is also solvable. By (ii) and $t_{0}^{\prime}\in\mathcal{T}$, we can apply the induction hypothesis to obtain $\mathcal{L}(t_{0}^{\prime})=\mathcal{L}(t_{2})$, and by (iii), we obtain $\mathcal{L}(t_{0})=\mathcal{L}(t_{1})$ as required. ∎
A.1.7 Proof of the Existence of Characteristic Data (Theorem 3.5)
@@
To prove Theorem 3.5, we first define characteristic learning data $M^{t}$ for each solvable tuple pattern $t$. Let $t=(p_{1},\ldots,p_{n})$ be a solvable tuple pattern. Since $t$ is solvable, $t$ contains at most $n$ distinct variables $x_{1},\ldots,x_{k}$ with $k\leq n$. (Note that, at each reduction step by $\leadsto$, the set of variables occurring in the tuple pattern does not change, and the number of elements may only monotonically decrease.) Let $a,b$ be two distinct variables of $\Sigma$, and $\alpha_{1},\ldots,\alpha_{n}$ be binary codewords for $1,\ldots,n$ using $a,b$ such that $|\alpha_{i}|=O(\log n)$ and for all $i,j(i\neq j)$, $\alpha_{i}$ is neither a prefix nor a suffix of $\alpha_{j}$. Let $\beta_{i}\,(i=1,\ldots,n)$ be $a/b,b/a]\alpha_{i}$. We define $M^{t}$ as: [^11] The characteristic data $M^{t}$ depends on the choice of $a,b\in\Sigma$ and binary prefix coding, but the choice does not matter for the discussion below.
@@
We use the following properties of characteristic data.
@@
  • Cases PR-CPrefix, PR-Suffix, and PR-CSuffix: Similar to the above case.
  • Case PR-Epsilon: We may assume without loss of generality that $t_{1}=(\epsilon,q_{2},\ldots,q_{n})$ and $t_{1}^{\prime}=(q_{2},\ldots,q_{n})$. By the assumption $M^{t},\Theta\models t_{1}$, $M^{t}[*][1]=\widetilde{\epsilon}$. Therefore, $t$ must be of the form $(\epsilon,p_{2},\ldots,p_{n})$, with $M^{t^{\prime}},\Theta\models t_{1}^{\prime}$ for $t^{\prime}=(p_{2},\ldots,p_{n})$. By the induction hypothesis, we have $\mathcal{L}(t^{\prime})\subseteq\mathcal{L}(t_{1}^{\prime})$. Thus, we have $\mathcal{L}(t)\subseteq\mathcal{L}(t_{1})$ as required.